# Data Whisperer: Efficient Data Selection for Task-Specific LLM Fine-Tuning via Few-Shot In-Context Learning

### v1

@@ 1 Introduction
Fine-tuning on task-specific data has become a standard approach for adapting large language models (LLMs) to specialized tasks (Ouyang et al., 2022). For instance, continued pre-training (Gururangan et al., 2020; Chang et al., 2024) involves extending the original pre-training phase on datasets closely aligned with the target domain; instruction tuning (Zhang et al., 2023; Üstün et al., 2024) focuses on leveraging instruction-response pairs to improve the model’s ability to follow instructions; and task-specific fine-tuning enhances model performance on a particular task by using domain-specific data (Yang et al., 2024b; Zhang et al., 2024). As datasets continue to expand in size, however, a critical challenge arises: how to efficiently select the optimal training examples—a process known as data selection—to strike a balance between computational cost and model performance (Zheng et al., 2023; Paul et al., 2021; Li et al., 2023b; Zhang et al., 2024; Li et al., 2023a).
@@ 2.2 Data Selection
Data selection is critical in deep learning, aiming to identify high-quality data that enhances training efficiency while maintaining comparable (or nearly lossless) performance. Traditional methods rely on heuristic metrics, with two primary strategies: (I) importance-based selection, which targets challenging or essential samples (Raffel et al., 2020; Paul et al., 2021; Mirzasoleiman et al., 2020; Killamsetty et al., 2021; Tan et al., 2024), and (II) similarity- or diversity-based selection, which selects samples based on their similarity to others or their representativeness of the feature space (Bukharin and Zhao, 2023; Xia et al., 2024; Zheng et al., 2023; Wei et al., 2021; Yu et al., 2024).
For LLMs, these methods face two challenges: (I) the unreliability of heuristic metrics in high-dimensional spaces, and (II) the computational cost of fine-tuning additional scoring models. In contrast, our approach directly integrates data selection with the pre-trained model’s predictions using ICL, eliminating the need for fine-tuning or heuristics.
@@ 3.4 Theoretical Analysis
$\mathcal{M}_{f}(q)$ $=(W_{V}+\Delta W_{V})X_{q,<t}$ (10)
$\cdot X_{q,<t}^{\top}(W_{K}+\Delta W_{K})^{\top}q$
$=\underbrace{W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q}_{\text{zero shot input}}$
$+\Delta W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q$
$+\Delta W_{V}X_{q,<t}(\Delta W_{K}X_{q,<t})^{\top}q$
$+\Delta W_{V}X_{q,<t}(\Delta W_{K}X_{q,<t})^{\top}q$
$=(W_{\textrm{zsl}}+\Delta W_{\textrm{ft}})q.$
@@ 4.3 Ablation Study
$\bullet$ Ablation of Attention Layers. The effectiveness of the context-aware weighting mechanism depends on the selected attention layer for scoring. We conducted an ablation study on the Llama-3-8B-Instruct model, evaluating shallow (Layer 5), intermediate (Layer 13), and deep (Layer 30) layers. As shown in Table 5, the intermediate layer (e.g., Layer 13) generally yielded more stable results, though all layers outperformed random selection. This suggests that Data Whisperer benefits from multiple attention layers, with intermediate layers providing more relevant contextual information for task-specific fine-tuning. These findings align with previous research showing that intermediate layers are crucial for semantic interpretation, processing moderately complex concepts, and providing contextual information for tasks requiring understanding (Wendler et al., 2024; Jin et al., 2025; Li et al., 2024).

### v3

@@ 1 Introduction
Fine-tuning on task-specific data has become a standard approach for adapting large language models (LLMs) to specialized tasks (Ouyang et al., 2022). For instance, continued pre-training (Gururangan et al., 2020; Chang et al., 2024) involves extending the original pre-training phase on datasets closely aligned with the target domain; instruction tuning (Zhang et al., 2023; Üstün et al., 2024) focuses on leveraging instruction-response pairs to improve the model’s ability to follow instructions; and task-specific fine-tuning enhances model performance on a particular task by using domain-specific data (Yang et al., 2024c; Zhang et al., 2024). As datasets continue to expand in size, however, a critical challenge arises: how to efficiently select the optimal training examples—a process known as data selection—to strike a balance between computational cost and model performance (Zheng et al., 2023; Paul et al., 2021; Li et al., 2023b; Zhang et al., 2024; Li et al., 2023a).
@@ 2.2 Data Selection
Data selection is critical in deep learning, aiming to identify high-quality data that enhances training efficiency while maintaining comparable (or nearly lossless) performance. Traditional methods rely on heuristic metrics, with two primary strategies: (I) importance-based selection, which targets challenging or essential samples (Raffel et al., 2020; Paul et al., 2021; Mirzasoleiman et al., 2020; Killamsetty et al., 2021; Tan et al., 2024), and (II) similarity- or diversity-based selection, which selects samples based on their similarity to others or their representativeness of the feature space (Bukharin and Zhao, 2023; Xia et al., 2024; Zheng et al., 2023; Wei et al., 2021; Yu et al., 2024; Yang et al., 2024b; Yu et al., 2025). For LLMs, these methods face two challenges: (I) the unreliability of heuristic metrics in high-dimensional spaces, and (II) the computational cost of fine-tuning additional scoring models. In contrast, our approach directly integrates data selection with the pre-trained model’s predictions using ICL, eliminating the need for fine-tuning.
2.3 Data Synthesis
Data synthesis generates additional high-quality data to enhance model performance and generalization. In LLMs, pretrained models are prompted to produce diverse, context-aware samples that augment scarce or imbalanced datasets (Wang et al., 2022; Li et al., 2024a; Xu et al., 2024), reducing annotation needs and enabling domain adaptation through iterative self-refinement. In vision, data distillation compresses large datasets or models into smaller, representative subsets or synthetic exemplars (Wang et al., 2020; Zhao et al., 2021; Cazenavette et al., 2022; Wang et al., 2025c, a, b), cutting redundancy and training costs while maintaining performance for efficient learning.
@@ 3.4 Theoretical Analysis
$\mathcal{M}_{f}(q)$ $=(W_{V}+\Delta W_{V})X_{q,<t}$ (10)
$\cdot X_{q,<t}^{\top}(W_{K}+\Delta W_{K})^{\top}q$
$=\underbrace{W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q}_{\text{zero shot input}}$
$+\Delta W_{V}X_{q,<t}(W_{K}X_{q,<t})^{\top}q$
$+\Delta W_{V}X_{q,<t}(\Delta W_{K}X_{q,<t})^{\top}q$
$+\Delta W_{V}X_{q,<t}(\Delta W_{K}X_{q,<t})^{\top}q$
$=(W_{\textrm{zsl}}+\Delta W_{\textrm{ft}})q.$
@@ 4.3 Ablation Study
$\bullet$ Ablation of Attention Layers. The effectiveness of the context-aware weighting mechanism depends on the selected attention layer for scoring. We conducted an ablation study on the Llama-3-8B-Instruct model, evaluating shallow (Layer 5), intermediate (Layer 13), and deep (Layer 30) layers. As shown in Table 5, the intermediate layer (e.g., Layer 13) generally yielded more stable results, though all layers outperformed random selection. This suggests that Data Whisperer benefits from multiple attention layers, with intermediate layers providing more relevant contextual information for task-specific fine-tuning. These findings align with previous research showing that intermediate layers are crucial for semantic interpretation, processing moderately complex concepts, and providing contextual information for tasks requiring understanding (Wendler et al., 2024; Jin et al., 2025; Li et al., 2024b).